OpenStudy (anonymous):

Graph the linear equation by solving For y first 2x+3y=3

OpenStudy (anonymous):

First get Y one side and others the other side

OpenStudy (anonymous):

y=1-2x/3

OpenStudy (anonymous):

now how to graph it

OpenStudy (anonymous):

then substitute X to -1, 1, 2, 3

OpenStudy (anonymous):

whatever you get Y and the numbers you substituted to X plot a graph

OpenStudy (anonymous):

2x + 3y = 3 3y = -2x + 3 y = -2/3x + 1 put a circle on the y line at 1. The slope is -2/3, so from the 1 on the y line, come down 2, and to the right 3, then down 2 and to the right 3. The x intercept is : 2x + 3y = 3.....2x + 3(0) = 3....2x = 3....x = 3/2 or 1 1/2. This is where the line crosses the x axis. It crosses the y axis at 1.

OpenStudy (anonymous):

i marked it wrong so i got it wrong i was off by a smudge

OpenStudy (anonymous):

i have a new one ugh i hate graphing

OpenStudy (anonymous):

x+5y=15

OpenStudy (anonymous):

to find the x intercept, sub 0 in for y to find the y intercept, sub in 0 for x to find the slope, put your equation in y = mx + b where m is your slope x + 5y = 15 5y = -x + 15 y = -1/5 + 3 3 is your y intercept. slope is -1/5....so from the 3 on your y intercept, come down 1(because it is negative) and to the right 5 (because it is positive) Your line should cross the x axis at : x + 5y = 15 x + 5(0) = 15 x = 15 it should cross the x axis at 15, and the y axis at 3

OpenStudy (anonymous):

or just go here....punch in graphing calculator....on the first line, write your equation, and go to the bottom and click graph... http://www.meta-calculator.com/online/